OpenStudy (anonymous):

can any1 help me with physics? i just need help with convertign stuff etc before putting stuff into the formula.

OpenStudy (anonymous):

There is a physics section. You can post there if you wish, i will be there also. But if you really insist on posting here, I don't mind a bit. Go ahead.

OpenStudy (anonymous):

oh thank you:), like some problems ays " a mass of 320 gm" then it has convertet it to 320 x 10^-3 ??

OpenStudy (anonymous):

320 x 10^-3 kg

OpenStudy (anonymous):

i know like problems are kind of easy, but the converting stuff is kinda hard :S

OpenStudy (anonymous):

gm stands for gram. 1000 gram is equivalent to 1 kilogram (kg). Thus 320 gram is equals to 320 x10 ^-3

OpenStudy (anonymous):

oh so the ^-3 is the amount of zeros?

OpenStudy (anonymous):

One trick is to count the decimal places lets say 1 x 10^ 3 That means shifting the decimal place 3 position to the right. Getting 1000. Lets say 1 x 10^-3. That means shifting the decimal place 3 position to the left. Getting 0.001.

OpenStudy (anonymous):

ohh i see :) thanks

OpenStudy (anonymous):

in some problems there are like; f=330/1.5 then from nowere it will be 3300/15 , why did they multiply a 10? so as to make it from kg to cm?

OpenStudy (anonymous):

In the example u mention. f= 330/1.5 and 3300/15. They are both the same because if you multiply 10 to both the numerator and denominator of 330/1.5, you will get 3300/15

OpenStudy (anonymous):

They are doing that for no particular purpose, the way i see it. It makes mental calculation easier? To me it is the same when you punch it all in the calculator.

OpenStudy (anonymous):

ye ur right thatnks for ur help :)

OpenStudy (anonymous):

and in some problems there is like ; v=wavelenght x frequency, then the question says " what is the wavelenght of the string" , should i solve for wavelenght making it ; wavelengt = v/f ??

OpenStudy (anonymous):

3.2x10^-2... I think he is talking about scientific notation